OpenStudy (anonymous):

Verify the identity. (1 - sinx)/cosx = cosx/(1 + sinx)

hartnn (hartnn):

hi @IBhardcore :) for your identity, what you can do as a first step is multiply and divide by 1+sin x \(\large \dfrac{1-\sin x}{\cos x} \: \dfrac{1+\sin x}{1+\sin x}\) now can you simplify the numerator ?

OpenStudy (anonymous):

sin ^2x + 1?

hartnn (hartnn):

you sure ? hint : \((a+b)(a-b) = a^2-b^2 \\ (1-\sin x )(1+\sin x)=...?\)

OpenStudy (anonymous):

1 - sin^2x

hartnn (hartnn):

correct! and using \(\sin^2 x+\cos^2 x =1\) what does 1 - sin^2x simplify to ???

OpenStudy (anonymous):

cos^2x

hartnn (hartnn):

good! now numerator has cos^2 x = cos x * cos x denominator has cos x (1+ sin x) notice that one cos x gets cancelled! what remains ?

OpenStudy (anonymous):

cos x / 1+ sin x !

hartnn (hartnn):

good! you got it :)

OpenStudy (anonymous):

Thanks!! :)

hartnn (hartnn):

welcome ^_^